[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Jul 20, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Jul 21 09:12:48 UTC 2017


HSU Shift Log for Thu Jul 20 2017 16:31:36 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: Wrote Josh Cossuth on the missing GMI data in the  
FNMOC/NRL pages. They have a permission issue and they are working on  
fixing it.

Miscellaneous issues: 00Z EC has a baroclinically generated TC from a MCV  
in the northern GOM day 7. Interestingly, the 12Z GFS gets close at the  
same time, but keeps it over land.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170721/5821cce0/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list